A 5.347 g of hexane undergo combustion.

g H₂O generated?


1
Expert's answer
2022-01-05T07:03:02-0500

The balanced chemical equation for the combustion of hexane is:

"2C_6H_{14}+19O_2 \to 12CO_2+14H_2O"

RMM for hexane = "6\times 12+1\times 14=86"

Moles of hexane used "=0.0623moles"

mole ratio "C_6H_{14}:CO_2=2:12"

Moles of "CO_2=12/2\times 0.0623=0.3738moles"

Mass of "CO_2=0.3738 \times (12+16+16)=16.4472g"

mole ratio "C_6H_{14}:H_2O=2:14"

Moles of "H_2O =14/2\times 0.0623=0.4361moles"

mass of "H_2O" produced = "0.4361\times (1+1+16)=7.8498"